OpenStudy (vishweshshrimali5):

Interesting advanced calculus question [\(\color{green}{\text{SOLVED}}\)]

OpenStudy (vishweshshrimali5):

\[\large{\color{red}{\text{QUESTION}}}\] For b > 0, prove that: 1. \(\large{\cfrac{\sin x}{x}dx = \int_{0}^{\infty}(\int_{0}^{b}e^{-ux}\sin x \ dx) \ du}\) 2. \(\large{|\int_{0}^{b}\cfrac{\sin x}{x}dx - \cfrac{\pi}{2}| < \cfrac{1}{b}}\)

OpenStudy (vishweshshrimali5):

\[\large{\color{red}{\text{HINT}}}\] For y > 0 prove that: \[\large{\int_{0}^{y}(\int_{0}^{b}e^{-ux}\sin x \ dx) \ du = \int_{0}^{b} (1-e^{-xy})\cfrac{\sin x}{x}dx}\] and then show that: \[\large{\lim_{y \rightarrow \infty}\int_{0}^{b} (1-e^{-xy})\cfrac{\sin x}{x}dx = \int_{0}^{b} \cfrac{\sin x}{x}dx}\]

OpenStudy (vishweshshrimali5):

\[\large{\color{red}{\text{SOLUTION}}}\] First of all we will prove that for b > 0 \[\large{\lim_{y \rightarrow \infty}\int_{0}^{b} (1-e^{-xy})\cfrac{\sin x}{x}dx = \int_{0}^{b} \cfrac{\sin x}{x}dx}\tag{1}\] We have for y > 0: \[\large{|\int_{0}^{b}\cfrac{\sin x}{x}dx - \int_{0}^{b}(1-e^{-xy})\cfrac{\sin x}{x}dx|}\] \[\large{=|\int_{0}^{b}e^{-xy}\cfrac{\sin x}{x}dx|}\] \[\large{\le \text{max}_{0 \le x \le b} |\cfrac{\sin x}{x}|\int_{0}^{b}e^{-xy} dx}\] \[\large{= M(b) \ [\cfrac{e^{-xy}}{-y}]_{0}^{b}}\] \[\large{= M(b) \ \cfrac{(1-e^{-by})}{y}}\] \[\large{\le \cfrac{M(b)}{y}}\] Letting \(\large{y \rightarrow \infty}\) we obtain equation 1... \(\large{\color{red}{;)}}\) Next we have: \[\large{\int_{0}^{b} (1-e^{-xy})\cfrac{\sin x}{x} dx}\] \[\large{=\int_{0}^{b}(\int_{0}^{y} xe^{-xu}du)\cfrac{\sin x}{x} dx}\] \[\large{=\int_{0}^{y}(\int_{0}^{b}e^{-ux}\sin x dx) du}\] Again letting \(\large{y \rightarrow \infty}\) we obtain: \[\large{\int_{0}^{b}\cfrac{\sin x}{x} dx = \int_{0}^{\infty}(\int_{0}^{b}e^{-ux}\sin x dx) du}\] \[\large{=\int_{0}^{\infty} \cfrac{(1-e^{-bu}(u\sin b + \cos b))}{1+u^2} du}\] \[\large{=\cfrac{\pi}{2} - \int_{0}^{\infty} \cfrac{e^{-bu}(u \sin b + \cos b)}{1+u^2} du}\] Thus we have: \[\large{|\int_{0}^{b}\cfrac{\sin x}{x} dx - \cfrac{\pi}{2}|}\] \[\large{=|\int_{0}^{\infty}\cfrac{e^{-bu}(u\sin b + \cos b)}{1+u^2} du|}\] \[\large{\le \int_{0}^{\infty}e^{-bu} \cfrac{\sqrt{1+u^2}}{1+u^2} \ du}\] \[\large{\le \int_{0}^{\infty}e^{-bu} du}\] \[\large{= \cfrac{1}{b}}\] \[\large{\color{red}{\text{Hence Proved}}}\] \[\large{\color{blue}{\text{- Prepared using Notepad++}}}\]

OpenStudy (vishweshshrimali5):

@Kainui @ganeshie8 Any other ideas/suggestion or anything :) ?

OpenStudy (vishweshshrimali5):

@ikram002p

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

What is this?

OpenStudy (vishweshshrimali5):

This is...uhm... @ganeshie8 what is this ?

OpenStudy (vishweshshrimali5):

This is \(\large{\color{red}{\text{MESS !!!!}}}\) @ℐℵK℣ⱱøƴȡ◆

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

Right =P

OpenStudy (vishweshshrimali5):

:D

OpenStudy (kainui):

I think I might have something for you give me a second! =)

OpenStudy (vishweshshrimali5):

Sure:)

ganeshie8 (ganeshie8):

i smell series/complex stuff from Kainui xD

OpenStudy (vishweshshrimali5):

He is making me nervous :( :P

OpenStudy (vishweshshrimali5):

And now he has stopped typing ... And now I am damn nervous :P :'(

OpenStudy (kainui):

Wait I messed up sorry guys to get you all riled up for nothing... X_X AHHHH!!! Sorry!!!

OpenStudy (vishweshshrimali5):

\[\color{red}{\text{ITS OKAYYYYYYYYYYY}}\] ;)

OpenStudy (vishweshshrimali5):

But what were you going to type @Kainui O.o

OpenStudy (kainui):

One second, I might have found my error.

OpenStudy (vishweshshrimali5):

kk

OpenStudy (ikram002p):

i loved this question :)

OpenStudy (kainui):

No, what I was thinking in my mind is exactly what you have done. I had thought there was another way that was similar which looked like this, but once I got closer I realized that I had remembered it incorrectly and your way was how I had seen it before. This is what I had typed: -- \[\Large I(t)=\int\limits_0^\infty \frac{\sin(tx)}{x}dx\] So we see that our problem is: \[\Large I(1)=\int\limits_0^\infty \frac{\sin(x)}{x}dx\] We differentiate I(t) with respect to t under the integral sign: \[\Large I'(t)=\int\limits_0^\infty \cos(tx)dx\] \[\Large I'(t)= \frac{1}{t}\sin(\infty)\] This is the same as \[\Large -1 \le I'(t) \le 1\] or \[\Large | \frac{1}{t} | \le 1\] Integrating with respect to t we have: \[\Large |\ln t +C | \le 1\] --- And here I stopped because something seemed wrong. X_X

OpenStudy (vishweshshrimali5):

Thanks @ikram002p :) @Kainui What went wrong ?

OpenStudy (ikram002p):

this is remineds me with fourier somehow ,,there is something like this mm (trying to remember )

OpenStudy (kainui):

I guess I could continue with this and get to a result worth something: \[\Large I(1)=\ln (1) + C = \int\limits_0^\infty \frac{\sin x}{x}dx\] \[\Large |\int\limits_0^\infty \frac{\sin x}{x}dx | \le 1\] I can't really see how to get anywhere further than this.

OpenStudy (ikram002p):

mmm can we use limits xd

OpenStudy (vishweshshrimali5):

Well we were using limits @ikram002p :P

OpenStudy (ikram002p):

hehe no other way than u did xD

OpenStudy (vishweshshrimali5):

:D LOL you would have seen that in my previous question ;)

OpenStudy (ikram002p):

well no im talking about something else , ill write down in my notes ,and ill post if it went correct :)

OpenStudy (vishweshshrimali5):

Good idea :)

OpenStudy (ikram002p):

hehe what i did is convert to comlex integral of e^i z /z but i couldnt rid of i even after finding residues when =0 mmm xD

OpenStudy (kainui):

Yeah I think I did something similar to you @ikram002p and it didn't work out haha. I separated it out and tried to make it into partial fractions and they ended up having complex coefficients. I guess that doesn't work.

OpenStudy (vishweshshrimali5):

Nothing wrong in trying though :) @ikram002p I am looking for some more possible methods as that will help me in having alternatives. Having another plan is always a good idea :D

OpenStudy (ikram002p):

yep , @Kainui i think its cuz complex integral help when we have a,b for integral mm so we could rid of complex coefficients

OpenStudy (vishweshshrimali5):

That would require some more work and then some more and more and some more ;)

OpenStudy (kainui):

I have found some interesting things with power series just now while looking at this. I will continue to see what I can find and maybe we'll have something fun to look at. =)

OpenStudy (vishweshshrimali5):

Great!!!!

OpenStudy (solomonzelman):

Your integrals look a little weird latex-wise, you have the upper and lower sums not on top and below the integral sign. I would suggest for a number `1.` or `2.` (looking at your first reply) to write the integrals your are having in an equation editor , and in front of the integral type `1.` and after it type `~` for space, you can type `~` more times for more space. On this site you won't get nice looking integrals or sigmas without an equation editor.

OpenStudy (vishweshshrimali5):

Thanks a lot @SolomonZelman . I was trying to figure out a way to do it since a long time. Thanks

OpenStudy (solomonzelman):

Anytime ... you can add before your integrals in this case \color{rsvssd}{ `1.~~~ ` } (I grayed the text you type And you can see it is without a specific color. It will be the closest to the numbers that look like the ones that are without the equation editor. No color, or some meaningless letters, or when you write " s `green` " (green but messed up (with an `s` or other not needed letter) in the color code space will give a gray-ish color

OpenStudy (solomonzelman):

or you can type them without an equation editor.

OpenStudy (solomonzelman):

The integrals I mean

OpenStudy (vishweshshrimali5):

Okay I will try to use them :)

OpenStudy (solomonzelman):

You can do `\(\large\color{blue}{ ^{ upper sum } }\)` and `\(\large\color{blue}{ _{ lower sum } }\)` but the sums will be a bit shifted though.

OpenStudy (solomonzelman):

Whatever you prefer, good luck and yw :)

OpenStudy (vishweshshrimali5):

:)